Can someone explain how to do this? I keep getting the wrong answer, I'm clearly not eliminating right?

J.C.​ Howard's medical testing company in Kansas wishes to assign a set of jobs to a set of machines. The following table provides the production data of each machine when performing the specific​ job:

Machine

​a) The optimal assignment of jobs to machine that will maximize total production​ is:

Machine A --

Machine B --

Machine C --

Machine D --

b)The total production of assignment = __ units (as a whole number)
Answer #1

(a)

Given data:

Job

Machine

A

B

C

D

1

9

11

5

8

2

12

8

13

9

3

7

9

8

10

4

10

9

7

6

The maximum production is 13

We have to substract each production data from 13

Job

Machine

A

B

C

D

1

13-9=4

13-11=2

13-5=8

13-8=5

2

13-12=1

13-8=5

13-13=0

13-9=4

3

13-7=6

13-9=4

13-8=5

13-10=3

4

13-10=3

13-9=4

13-7=6

13-6=7

Final table will look like as follows.

Job

Machine

A

B

C

D

1

4

2

8

5

2

1

5

0

4

3

6

4

5

3

4

3

4

6

7

Identify the minimum for each row.

Job

Machine

A

B

C

D

1

4

2

8

5

2

1

5

0

4

3

6

4

5

3

4

3

4

6

7

(a)

Machine A: Job 4

Machine B: Job 1

Machine C: Job 2

Machine D: Job 3

(b)

Total production:

10+11+13+10 = 44